# Demonstration that Differential Length Changes of Optical Cavities are a Sensitive Probe for Ultralight Dark Matter

### v2

@@ Abstract
Measurements of differential length oscillations of Fabry-Perot cavities provide a sensitive and promising approach to searching for scalar ultralight dark matter (ULDM). The initial demonstration sets direct lower bounds that are one to two orders of magnitude lower for two model ULDM distributions – a standard galactic halo, and a relaxion star bound to Earth – ranging over a decade of ULDM mass and Compton frequency. The demonstration suggests how a much higher sensitivity to a much larger ULDM mass range can be obtained.
@@
This Letter reports a new approach to discovering or obtaining bounds on a well-motivated alternative – scalar, parity-even, ultralight dark matter (ULDM) that behaves as a classical wave [7]. Fig. 1 represents a demonstration apparatus that establishes new bounds that are lower by one to two orders of magnitude, over a tenfold change in ULDM mass and Compton frequency. Also indicated are paths to a much higher sensitivity over a much larger mass range, and the feasibility of operating at cryogenic temperatures to reduce the effect of thermal noise as this becomes increasingly important. The measurement detects differing length changes of two cryogenic cavities at the ULDM Compton frequency $f_{\phi}\equiv m_{\phi}c^{2}/(2\pi\hbar)$, for a ULDM mass $m_{\phi}$, Planck constant $\hbar$, and speed of light $c$. Sensitivity to the local dark matter density produces strong and direct limits on boson stars [8, 9] composed of relaxions – a proposed solution to the hierarchy problem [10] – a contrast to the indirect limits from torsion balance equivalence principle (EP) tests [11, 12].
ULDM is well-motivated, being consistent with the standard halo model (SHM) distribution of dark matter in a galaxy [12], and naturally occurring in BSM (beyond the SM) theories. The topologically complex vacua of string theory, for example, produce many ULDM candidates. These include moduli and dilatons [13, 14, 15, 16, 17, 18, 19] that couple to matter by making SM parameters depend on moduli fields. The moduli can acquire mass through supersymmetry (SUSY) breaking at $\sim 0.1\,{\rm meV}$ for TeV scale SUSY [13], or have much lighter masses due to loop factors and small coefficients (e.g. for the electron Yukawa modulus). The electron mass and charge ($m_{\rm e}$ and ($-e$), and the fine structure constant ($\alpha$), vary in time ($t$) and space ($\mathbf{x}$) in proportion to the ULDM field $\phi(t,\mathbf{x})$,
@@
The ULDM coherence length, $\lambda_{\phi}\equiv\hbar/(m_{\phi}v_{{\rm vir}})$ [21] is much larger than our tabletop apparatus for both ULDM models, so $\phi(t,\mathbf{x})\approx\phi(t)$ in Eq. (1). The virial velocity $v_{{\rm vir}}$ is for ULDM in equilibrium, with mean kinetic energy equal to $(-1/2)$ multiplied by the mean gravitational potential energy. For the SHM, $v_{\rm vir}\approx 166\,{\rm km}/{\rm s}$ ([4] and Appendix D) so $\lambda_{\phi}$ ranges from 8 ×$\times 10^{2}$ to 172 ×$\times 10^{2}$ km for the $f_{\phi}$ between 5 and 100 kHz on which we set new bounds. For a relaxion star gravitationally bound to the Earth, $v_{\rm vir}\approx 32\,{\rm m}/{\rm s}\times(f_{\phi}/1\,{\rm kHz})$ [10] so $\lambda_{\phi}$ ranges from $5\times 10^{4}$ to $110\times 10^{4}\,{\rm km}$ for the $20$ to 909090\,90kHz frequency range we consider.
@@
The ULDM coherence times $\tau_{{\rm c}}\equiv\hbar/(m_{\phi}v_{\rm vir}^{2})$ [21, 26] can range from short ($\tau_{\rm c}\ll T_{\rm m}$) to long ($\tau_{\rm c}\gg T_{\rm m}$) compared to our measurement time, $T_{\rm m}\approx 4\,\rm{days}\approx 4\times 10^{5}\,\rm{s}.$ For the relaxion model, the coherence times for the range of $f_{\phi}$ we consider go from short ($\tau_{\rm c}=2\times 10^{4}\,\rm{s}$) to long ($\tau_{\rm c}=2\times 10^{6}\,\rm{s}$). For the SHM, the coherence time is always short since $\tau_{c}$ ranges only from $5\,\rm{s}$ to $10^{2}\,\rm{s}$ over the range of $f_{\phi}$ we consider.
For long coherence times, $\phi(t)\approx\Phi_{0}\cos(2\pi f_{\phi}t+\theta)$ with a stochastic phase $\theta$, and amplitude $\Phi_{0}\sim\sqrt{\rho_{{\scriptscriptstyle{\rm DM}}}}$, that goes as the square root of the local dark matter density. For short coherence times, a sum over many Fourier components is required, in a frequency window around $f_{\phi}$ with a width that scales as $\tau_{\rm c}^{-1}$. All components go as $\sqrt{\rho_{{\scriptscriptstyle{\rm DM}}}}$. One result is the so-called “ULDM lineshape” [21], an example of which is shown for the SHM in Fig. 2 (a). All coherence times are appropriately accounted for in our Bayesian data analysis ([21, 26] and Appendix D).
@@
Both of the cryogenic optical cavities are made of single-crystal sapphire due to its high thermal conductivity, large Young’s modulus, and low intrinsic loss at cryogenic temperatures [27]. Their longitudinal axes are parallel to the crystallographic c-axis to maximize longitudinal stiffness. Optically contacted sapphire mirrors are coated with low-Brownian-noise crystalline GaAs/AlGaAs dielectric Bragg reflectors [28].
Fig. 1 (b) is a overview of the cavities and laser system. The amplified light from a fiber laser with frequency $f_{{\rm opt}}\approx 193\,{\rm THz}$ is Pound-Drever-Hall (PDH) locked [29] to a stable, ultralow expansion (ULE) cavity that is 10 cm long with a finesse of 500,000. The light is then split into beams frequency shifted by $\approx 131\,{\rm MHz}$ (Fig. 6) to make them nearly resonant with cavities that are $L_{\scriptscriptstyle{{\rm L}}}=15\,{\rm cm}$ and $L_{\scriptscriptstyle{{\rm S}}}=7.5\,{\rm cm}$ long. The differential “optical strain” for fluctuations $\delta f_{{\rm beat}}(t)$ in their 262 MHz beat note is
@@
The ULDM $f_{\phi}$ primarily searched for are well above the 4.7 and 7.6 kHz optical poles for the long and short cavites. The detector response function of Eq. (3) then simplifies (Appendix C) to the difference
@@
Each has its own resonant frequency $f_{{\scriptscriptstyle{\rm M}}}$ and large quality factor $Q_{{\scriptscriptstyle{\rm M}}}$. The second subscript in Eq. (4) designates the long (L) or short (S) cavity. Each sapphire cavity acts on a ULDM signal as a low pass filter, given that $Q_{{\scriptscriptstyle{\rm M}}}^{-1}$ is negligible away from $f_{M}$. For frequencies above $f_{{\scriptscriptstyle{\rm M}}}$, the cavity does not follow the ULDM-induced oscillations [24].
Figure 3: Qualitative magnitude and phases of two cavities’ individual mechanical transfer functions (Eq. (5)), $H_{{\scriptscriptstyle{\rm M,\,L/S}}}$ for the long/short cavity, and the detector response function $A_{{\rm det}}$ (Eq. (4)) with the primary science band highlighted in gray.
Fig. 3 shows magnitudes and phases of low pass transfer functions representing long and short cavities, whose difference determines $A_{\rm det}$. The highest sensitivity to $h_{{\scriptscriptstyle{\rm DM}}}$ is in our “science band” between the longitudinal mechanical resonances of the two cavities at $f_{{\scriptscriptstyle{\rm M\,,L}}}\approx 34\,$kHz and $f_{{\scriptscriptstyle{\rm M\,,S}}}\approx 64\,$kHz. The short cavity approximately follows the ULDM drive, while the long cavity (largely unaffected by ULDM) is a length reference.
@@
The ASD of the large data set $\widetilde{h}(f)$ is analyzed to look for evidence of the two different ULDM models using Bayesian analysis [21, 26]. In a frequency window as wide as the ULDM lineshape (e.g. Fig. 2 (a)), we compute the likelihood that a ULDM signal $\widetilde{s}_{{\scriptscriptstyle{\rm DM}}}$ (Fourier transform of $A_{{\rm det}}h_{{\scriptscriptstyle{\rm DM}}}(t)$ for a given $d_{m_{{\rm e}}}$) gives rise to the measured data $\widetilde{h}(f)$. We use the likelihood function, ($\mathcal{L}(\{\widetilde{h}\}|d_{m_{{\rm e}}})$, defined in Ref. [21, 26]. The probability distribution function of the unknown parameter $d_{m_{\rm e}}$, known as the posterior, $\mathcal{P}(d_{m_{{\rm e}}}|\{\widetilde{h}\})$, is computed from the likelihood using Bayes theorem (Appendix D). Solving
@@
gives $d_{m_{\rm e}}$ at a 95% confidence level.
In Fig. 4 (a), our increased sensitivity (gray area) and much smaller ULDM limits for the SHM (solid black curve) come from solving Eq. (6). This is for a terrestrial detector moving through the galactic dark matter halo towards the Cygnus constellation, as the Solar system orbits the Milky Way, with a relative mean speed of $v_{\rm obs}\approx 230\,{\rm km}/{\rm s}$ and a Gaussian velocity spread with $v_{\rm vir}\approx 166\,{\rm km}/{\rm s}$ [4] (Appendix D). We look for spectral peaks in $\widetilde{h}(f)$ whose profile and 0.01 to 0.2 Hz widths are set by the ULDM lineshape and inverse coherence time (Fig. 2 (a)).
@@
Fig. 5 displays our SHM result for a much broader range of coupling constants and Compton frequencies to relate the new bounds to other measurement bounds and theoretical constraints. In addition to those shown in Fig. 4 (a) and (b), these include direct ULDM limits from gravitational wave (GW) detectors [33, 34] and the Holometer [35], alongside indirect bounds from EP tests [12, 36, 37]. Radiative corrections to the ULDM mass provide a very model-dependent upper limit on $d_{m_{\rm e}}$, with the parameter space above the dashed line from [12] labeled “naturalness” excluded without substantial fine-tuning [24]. In addition, we set lower limits in the 6 to 10 kHz range for the SHM, where the previous strongest limits came from molecular Iodine spectroscopy [38].
@@
Figure 5: Bounds on $d_{m_{\rm e}}$ for the SHM for a much larger range of $d_{m_{\rm e}}$ and Compton frequencies compared to other bounds set by direct and indirect measurements and by theoretical limits. A zoom in of the dashed red box is shown in Fig. 4 (a). See main text for details.
The demonstration measurement, limited by detector noise that is ten times the photon shot noise, suggests the possibility of very large improvements. The sensitivity could be improved by 3 to 4 orders of magnitude using laser noise filtering and vibration isolation (in addition to what has been demonstrated), along with increased laser power to reduce the shot noise, whereupon the thermal noise would be reduced by the demonstrated cryogenic environment. The frequency and mass reach can be greatly extended using a longer science cavity and/or a non-rigid reference cavity, down to frequencies of 100 Hz [24]. The greatly improved broadband sensitivity could be used to search for ULDM from a boson star explosion in a “bosenova” near Earth [9] or from binary black hole and neutron star mergers [39].
Switching from our current broadband methods to resonant methods that use the quality factors of $10^{4}$ already observed could bring very large increases in sensitivity. More precise characterization of the mechanical resonance frequencies would allow a much deeper search in future experiments at these particular frequencies (Fig. 7) [40, 41]. Future measurements could even exploit the material properties of cryogenic sapphire cavities to realize quality factor values as high as $10^{8}$ [27, 42].
Two perpendicular cavities would give sensitivity to GWs above the 10 kHz upper limit of LIGO [43]. We estimate a GW strain sensitivity of $\sim 10^{-19}\,\text{Hz}^{-1/2}$ at the $34\,{\rm kHz}$ and $64\,{\rm kHz}$ mechanical resonances of our current cavities. With the mentioned sensitivity improvements, it could become possible to probe the theoretically well-motivated parameter space for GWs from primordial black hole mergers or gravitationally bound states of QCD axions [43, 44].
@@
In conclusion, a demonstration of a fresh approach to broadband direct ULDM detection sets new lower limits on the coupling of ultralight dark matter (ULDM) and ordinary matter. The new bounds are up to two orders of magnitude lower for both the standard halo model (SHM) and for a relaxion star bound to Earth. Routes to sensitivity increases of 3 to 4 orders of magnitude, for a much broader range of ULDM masses now seem feasible, with even greater sensitivities at the mechanical resonance frequencies of the optical cavities. Perpendicular cavities could provide gravitational wave sensitivity at frequencies above the 10 kHz limit of current large detectors.
@@
Acknowledgments: The Templeton Foundation funded the apparatus construction and the Packard Foundation funded T. K. B. Odom contributed to early stages of this work, and useful suggestions came from C. Schreibeis, A. Derevianko, G. Cole, S. Catano, R. DeSalvo, K. Arai, Y. Michimura, G. Venugopalan, and J. Bourhill. The custom dewar, inverted pendulum, and geometric anti-spring system were built by Four Nine Design to our specifications.

### v3

@@ Abstract
Measurements of differential length oscillations of Fabry-Perot cavities provide a sensitive and promising approach to searching for scalar ultralight dark matter (ULDM). The initial demonstration sets direct lower bounds that are one to two orders of magnitude lower for two model ULDM distributions – a standard galactic halo and a relaxion star bound to Earth – ranging over a decade of ULDM mass and Compton frequency. The demonstration suggests how a much higher sensitivity to a much larger ULDM mass range can be obtained.
@@
This Letter reports a new approach to discovering or obtaining bounds on a well-motivated alternative – scalar, parity-even, ultralight dark matter (ULDM) that behaves as a classical wave [7]. Fig. 1 represents a demonstration apparatus to search for ULDM by comparing the lengths of two cavities with different mechanical responses to ULDM. This approach offers superior sensitivity to length changes at the ULDM Compton frequency, $f_{\phi}\equiv m_{\phi}c^{2}/(2\pi\hbar)$, in the 5 kHz to 100 kHz range, outperforming competing methods [8, 9] by up to 2 orders of magnitude. Here, $m_{\phi}$, $\hbar$ and $c$ are the ULDM mass, Planck’s constant and the speed of light respectively. Moreover, unlike other cavity-based schemes exploring $f_{\phi}\gtrsim 1\,{\rm kHz}$ [8, 9], the demonstration of a vibrationally isolated cryogenic setup opens pathways to significant sensitivity enhancement (up to 6 orders of magnitude) over a broader $f_{\phi}$ range (1 kHz to 1 MHz), thus uniquely allowing exploration of theoretically motivated parameter space. Furthermore, sensitivity to the local dark matter density produces strong and direct limits on boson stars [10, 11] composed of relaxions – a proposed solution to the hierarchy problem [12] – a contrast to the indirect limits from torsion balance equivalence principle (EP) tests [13, 14].
ULDM is well-motivated, being consistent with the standard halo model (SHM) distribution of dark matter in a galaxy [14], and naturally occurring in BSM (beyond the SM) theories. The topologically complex vacua of string theory, for example, produce many ULDM candidates. These include moduli and dilatons [15, 16, 17, 18, 19, 20, 21] that couple to matter by making SM parameters depend on moduli fields. The moduli can acquire mass through supersymmetry (SUSY) breaking at $\sim 0.1\,{\rm meV}$ for TeV scale SUSY [15], or have much lighter masses due to loop factors and small coefficients (e.g. for the electron Yukawa modulus). The electron mass and charge ($m_{\rm e}$ and (e-e), and the fine structure constant ($\alpha$), vary in time ($t$) and space ($\mathbf{x}$) in proportion to the ULDM field $\phi(t,\mathbf{x})$,
@@
The ULDM coherence length, $\lambda_{\phi}\equiv\hbar/(m_{\phi}v_{{\rm vir}})$ [23] is much larger than our tabletop apparatus for both ULDM models, so $\phi(t,\mathbf{x})\approx\phi(t)$ in Eq. (1). The virial velocity $v_{{\rm vir}}$ is for ULDM in equilibrium, with mean kinetic energy equal to (1/2)(-1/2) multiplied by the mean gravitational potential energy. For the SHM, $v_{\rm vir}\approx 166\,{\rm km}/{\rm s}$ ([4] and Appendix E) so $\lambda_{\phi}$ ranges from 8 $\times 10^{2}$ to 172 $\times 10^{2}$ km for the $f_{\phi}$ between 5 and 100 kHz on which we set new bounds. For a relaxion star gravitationally bound to the Earth, $v_{\rm vir}\approx 32\,{\rm m}/{\rm s}\times(f_{\phi}/1\,{\rm kHz})$ [12] so $\lambda_{\phi}$ ranges from $5\times 10^{4}$ to $110\times 10^{4}\,{\rm km}$ for the $20$ to 9090\,kHz frequency range we consider.
@@
The ULDM coherence times $\tau_{{\rm c}}\equiv(c/v_{{\rm vir}})^{2}/(2\pi f_{\phi})$ [23, 28] can range from short ($\tau_{\rm c}\ll T_{\rm m}$) to long ($\tau_{\rm c}\gg T_{\rm m}$) compared to our measurement time, $T_{\rm m}\approx 4\,\rm{days}\approx 4\times 10^{5}\,\rm{s}.$ For the relaxion model, the coherence times for the range of $f_{\phi}$ we consider go from short ($\tau_{\rm c}=2\times 10^{4}\,\rm{s}$) to long ($\tau_{\rm c}=2\times 10^{6}\,\rm{s}$). For the SHM, the coherence time is always short since $\tau_{c}$ ranges only from $5\,\rm{s}$ to $10^{2}\,\rm{s}$ over the range of $f_{\phi}$ we consider.
For long coherence times, $\phi(t)\approx\Phi_{0}\cos(2\pi f_{\phi}t+\theta)$ with a stochastic phase $\theta$, and amplitude $\Phi_{0}\equiv c\sqrt{2\rho_{{\scriptscriptstyle{\rm DM}}}}/(2\pi f_{\phi})$, that goes as the square root of the local dark matter density. For short coherence times, a sum over many Fourier components is required, in a frequency window around $f_{\phi}$ with a width that scales as $\tau_{\rm c}^{-1}$. All components go as $\sqrt{\rho_{{\scriptscriptstyle{\rm DM}}}}$. One result is the so-called “ULDM lineshape” [23], an example of which is shown for the SHM in Fig. 2 (a). All coherence times are appropriately accounted for in our Bayesian data analysis ([23, 28] and Appendix E).
@@
Both of the cryogenic optical cavities are made of single-crystal sapphire due to its high thermal conductivity, large Young’s modulus ($\approx 400\,{\rm GPa}$), and low intrinsic loss at cryogenic temperatures [29]. Their longitudinal axes are parallel to the crystallographic c-axis to maximize longitudinal stiffness. Optically contacted sapphire mirrors are coated with low-Brownian-noise crystalline GaAs/AlGaAs dielectric Bragg reflectors with Young’s modulus and loss factor of $\approx 100\,{\rm GPa}$ and $\approx 5\times 10^{-6}$ respectively [30].
Fig. 1 (b) is an overview of the cavities and laser system. Amplified light from a fiber laser, with frequency $f_{{\rm opt}}\approx 193\,{\rm THz}$, is Pound-Drever-Hall locked to a $10\,{\rm cm}$, 500,000-finesse ultralow expansion (ULE) cavity. The light is then split and frequency-shifted by $\approx\pm 131\,{\rm MHz}$ (Fig. 6) to resonate with cavities of length $L_{\scriptscriptstyle{{\rm L}}}=15\,{\rm cm}$ and $L_{\scriptscriptstyle{{\rm S}}}=7.5\,{\rm cm}$ ($\approx 0.6\,{\rm mW}$ input per cavity). The frequency shift closes the gap in the cavities’ optical resonances. The differential “optical strain” for fluctuations $\delta f_{{\rm beat}}(t)$ in their 262 MHz beat note is
@@
Figure 3: (a) Qualitative magnitude and phases of two cavities’ individual mechanical transfer functions (Eq. (5)), $H_{{\scriptscriptstyle{\rm M,\,L/S}}}$ for the long/short cavity, and the detector response function $A_{{\rm det}}$ (Eq. (4)). (b) Dashed curves represent $A_{\rm det}$ for the lowest and highest data-driven estimates for the mechanical resonances of the two cavities with $Q_{{\scriptscriptstyle{\rm M}}}\sim 10^{4}$. The black curve conservatively takes their minimum to set ULDM limits.
The ULDM $f_{\phi}$ primarily searched for are well above the optical poles of the long and short cavities. The detector response function of Eq. (3) then simplifies (Appendix D) to the difference
@@
Each has its own resonant frequency $f_{{\scriptscriptstyle{\rm M}}}$ and quality factor $Q_{{\scriptscriptstyle{\rm M}}}$. The second subscript in Eq. (4) designates the long (L) or short (S) cavity. Each sapphire cavity acts on a ULDM signal as a low pass filter, given that $Q_{{\scriptscriptstyle{\rm M}}}^{-1}$ is negligible away from $f_{{\scriptscriptstyle{\rm M}}}$. Above $f_{{\scriptscriptstyle{\rm M}}}$, the cavity does not follow the ULDM-induced oscillations [26].
We developed a data-driven model of Eq. (4) by empirically estimating $f_{{\scriptscriptstyle{\rm M}}}$ and $Q_{{\scriptscriptstyle{\rm M}}}$ and fitting to finite-element analysis simulations. Fig. 3 qualitatively shows magnitudes and phases of low pass transfer functions representing long ($f_{{\scriptscriptstyle{\rm M\,,L}}}\approx 34\,$kHz) and short ($f_{{\scriptscriptstyle{\rm M\,,S}}}\approx 64\,$kHz) cavities, whose difference determines $A_{\rm det}$. The short cavity approximately follows the ULDM drive, while the long cavity (largely unaffected by ULDM) is a length reference. The ULDM limits shown in Fig. 4 & 5 account for the uncertainty in the model (Appendix D).
@@
[…]
@@
gives $d_{m_{\rm e}}$ at a 95% confidence level. The detector response function, $A_{{\rm det}}$, implicit in Eq. (6) ensures the noise floor does not determine $d_{m_{{\rm e}}}$ in the parameter regime where the laser beam does not carry the dark matter signal.
In Fig. 4 (a), our increased sensitivity (gray area) and much smaller ULDM limits for the SHM (solid black curve) come from solving Eq. (6). This is for a terrestrial detector moving through the galactic dark matter halo towards the Cygnus constellation, as the Solar system orbits the Milky Way, with a relative mean speed of $v_{\rm obs}\approx 230\,{\rm km}/{\rm s}$ and a Gaussian velocity spread with $v_{\rm vir}\approx 166\,{\rm km}/{\rm s}$ [4] (Appendix E). We look for spectral peaks in $\widetilde{h}(f)$ whose profile and 0.01 to 0.2 Hz widths are set by the ULDM lineshape and inverse coherence time (Fig. 2 (a)). For $T_{\rm m}\gg\tau_{{\rm c}}$, as is the case for the SHM, an approximate analytic solution to Eq. (6) is $d_{m_{{\rm e}}}\approx 2.56\,\widetilde{h}/(A_{{\rm det}}A_{0}\Phi_{0}(T_{\rm m}\,\tau_{{\rm c}})^{1/4})$, where $A_{0}\equiv\sqrt{4\pi\hbar c}/E_{{\scriptscriptstyle{\rm P}}}$ [23]. The $f_{\phi}^{-1}$ scaling of $\Phi_{0}$ and $\tau_{{\rm c}}$ and accounting for the optical poles in $A_{{\rm det}}$ (Appendix D) compensates for the increase in $\widetilde{h}$ (Fig. 2 (b)) at lower frequencies.
@@
Fig. 5 displays our SHM result for a much broader range of coupling constants and Compton frequencies to relate the new bounds to other measurement bounds and theoretical constraints. In addition to those shown in Fig. 4 (a) and (b), these include direct ULDM limits from gravitational wave (GW) detectors [32, 33] and the Holometer [34], along with indirect limits from EP tests [14, 35, 36]. Radiative corrections to the ULDM mass provide a very model-dependent upper limit on $d_{m_{\rm e}}$, with the parameter space above the dashed line from [14] labeled “naturalness” excluded without substantial fine-tuning [26]. In addition, we set lower limits in the 6 to 10 kHz range for the SHM, where the previous strongest limits came from molecular Iodine spectroscopy [37].
@@
Figure 5: Bounds for the SHM for a much larger range of $d_{m_{\rm e}}$ and Compton frequencies compared to other bounds set by direct and indirect measurements, our sensitivity projections and theoretical limits. A zoom in of the dashed red box is shown in Fig. 4 (a). See main text for details.
Furthermore, Fig. 5 shows theoretical $d_{m_{\rm e}}$ projections limited by the cavities’ thermal and shot noise for the current (“projection (now)”) and proposed future apparatuses (Appendix C). The latter is shown for two configurations $L_{\scriptscriptstyle{{\rm L}}}=1\,{\rm m},\,L_{\scriptscriptstyle{{\rm S}}}=30\,{\rm cm}$ (“projection (future 1)”) and $L_{\scriptscriptstyle{{\rm L}}}=1\,{\rm m},\,L_{\scriptscriptstyle{{\rm S}}}=1\,{\rm cm}$ (“projection (future 2)”). Projections assume temperature, quality factor ($\approx 10^{4}$), and finesse similar to our current apparatus (Appendix A). However, measurement time ($T_{\rm m}=10^{7}\,{\rm s}$) and input laser power (50 mW per cavity) were increased. Cavity lengths affect the explored parameter space; the first configuration prioritizes sensitivity over bandwidth, exploring below the naturalness line. We attribute the discrepancy in the measured and theoretical sensitivities to residual laser frequency/amplitude noise, which we plan to mitigate with additional optical filtering in future apparatuses.
Our projected sensitivities highlight the advantage of our method over other cavity-based ULDM detection schemes in Fig. 5. Schemes involving comparison of cavities to atoms [9] or molecules [37] are limited by factors like lower atom flux and broad atomic transition linewidths. Cavities benefit from a much higher photon flux, lower cavity linewidth and low thermal Brownian noise. While the cavity in Ref. [8] shares these benefits, their self-referential nature results in loss of sensitivity to ULDM above the cavity’s mechanical resonance. Moreover, fiber delay lines suffer from acoustic and thermal noise of the long fiber at low frequencies. Dual-cavity optical interferometers like GW detectors do not improve ULDM sensitivity with extended Fabry-Perot cavities. Since only beam splitter oscillations are monitored, their sensitivity is fundamentally comparable to their tabletop counterparts [27].
@@
In conclusion, a demonstration of a fresh approach to broadband direct ULDM detection sets new limits on the coupling of ultralight dark matter (ULDM) and ordinary matter. The new bounds are up to two orders of magnitude lower for both the standard halo model (SHM) and for a relaxion star bound to Earth. Routes to sensitivity increases to exceed the naturalness bound and a much broader range of ULDM masses now seem feasible.
@@
Acknowledgments: The Templeton Foundation funded the apparatus construction and the Packard Foundation funded T. K.; B. Odom contributed to early stages of this work and useful suggestions came from C. Schreibeis, A. Derevianko, G. Cole, S. Catano, R. DeSalvo, K. Arai, Y. Michimura, G. Venugopalan, and J. Bourhill. The custom dewar, inverted pendulum, and geometric anti-spring system were built by Four Nine Design to our specifications.